OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. A^1=-2 A^2=8 A^3=-32

OpenStudy (anonymous):

and its not An=-2*an-1

OpenStudy (anonymous):

well the first thing i see is it alternates between positive and negative. do you know how to write that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\huge (-1)^n\] do you see how we would jump back and forth between -1 and 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but on all of my answers it stays -1

OpenStudy (anonymous):

Great! now all we need to figure out is how it goes from 2, 8 and then 32.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Is this a multiple choice question or do you have an answer booklet of some sort?

OpenStudy (anonymous):

I could be misunderstanding the question. :/

OpenStudy (anonymous):

A. an=-4+an-1 B. an=-2+an-1 D. an=-4*an-1

OpenStudy (anonymous):

th reason i didnt write c is because its the wrong answer

OpenStudy (anonymous):

Ah, i see what we need to do here. http://www.algebralab.org/lessons/lesson.aspx?file=algebra_geoseq.xml We need to find the common ratio and then we'll have our answer. It talks about that in the first paragraph or so on that website.

OpenStudy (anonymous):

so we just need to know what we can multiply the first term by that gives us 8 and then that number has to work when multiplying it by 8 to get 32. Does that make sense?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So basically that pattern i was talking about before. Just ignore my (-1)^n stuff. That's for a different representation okay?

OpenStudy (anonymous):

So you can setup an equation to figure it out. :D \[\huge -2*x=8\]\[\huge 8*x=-32\] x must be the same number for both answer. give it a try.

OpenStudy (anonymous):

Once you find x, your answer will be in the form: \[\huge a_n=a_{n-1}*x\] I hope that all makes sense! :D